OpenStudy (_xoxo_hearts_):

Lila earns $8.75/h at her job. How many hours must Lila work to earn $105? Complete the equation to model the problem. Use h to represent the unknown value. @dan815 can you please help me?

OpenStudy (dan815):

let h be the number of hours lila works let M be the money she makes M=8.75*h she needs to make $105 105=8.75*h , solve for h

OpenStudy (_xoxo_hearts_):

so im working the problem out backwards?

OpenStudy (dan815):

This is how you set up an equatioon

OpenStudy (_xoxo_hearts_):

is it 12

OpenStudy (dan815):

yep

OpenStudy (dan815):

8.75*12 = 105

OpenStudy (_xoxo_hearts_):

thank you will you be online tomorrow I kinda need I tutor and I need to catch up big time and I have a end of the year test next week